b1az34:

The pay rate and hours worked are given below. Use this information to determine the following. Make sure to include decimals and appropriate zeros. the gross pay federal tax (assuming 20% of gross earnings) state tax (assuming 6% of gross earnings) social security deductions (assuming 7.05% of gross earnings) group health insurance (assuming 13% of gross earnings) pension deduction (assuming 6.5% of gross earnings) total deductions net pay Earnings Description Rate Hours Current Regular 9.00 39.0 351.00

b1az34:

KyledaGreat:

The net pay of the individual after having deducted total deductions is $166.54 How do we determine the amount of each deduction? The amount of each deduction is based on the gross pay of $351.00, as a result, we simply multiply the percentage of deduction by the gross pay to determine the dollar amount of each deduction. the gross pay=$351 federal tax (assuming 20% of gross earnings)=$351*20%=$70.20 state tax (assuming 6% of gross earnings)=$351*6%=$21.06 social security deductions (assuming 7.05% of gross earnings)=$351*7.05%=$24.75 group health insurance (assuming 13% of gross earnings)=$351*13%=$45.63 pension deduction (assuming 6.5% of gross earnings)=$351*6.5%=$22.82 total deductions=$70.20+$21.06+$24.75+$45.63+$22.82 total deductions=$184.46 net pay=$351-$184.46 net pay=$166.54